Please solve by completing the square
x^2 -6x = 27

Answer by mananth(16946)

x^2-6x=27
x^2-6x+9=27+9
(x-3)^2= 36
square root
(x-3)=+/-6
x= 9,-3
..
m.ananth@hotmail.ca